The population means and standard deviation are given below. Find the required probability and determine whether the given sample mean would be considered unusual.
For a sample of n equals 70, find the probability of a sample mean being less than 19.5 if u equals20 and o equals1.19.

Answer by stanbon(75887)

The population means and standard deviation are given below. Find the required probability and determine whether the given sample mean would be considered unusual.
For a sample of n equals 70, find the probability of a sample mean being less than 19.5 if u equals 20 and o equals 1.19.
-------
z(19.5) = (19.5-20)/(1.19/sqrt(70)) = -3.5154
P(x-bar < 19.5) = P(z < -3.5154) = normalcdf(-100,-3.5154) = 0.0002196
-----------
Cheers,
Stan H.
-------------